# Assessing the dynamical assumptions in Tsirelson inequality tests of non-classicality in harmonic oscillators

### v2

@@ Abstract
“Macrorealism” posits that a system possesses definite properties at all times and that we can discover these properties, in principle, without disturbing the system’s subsequent behaviour. The Leggett–Garg inequalities are derived under these assumptions and are readily violated by standard quantum mechanics, thereby providing a scheme to test whether demonstrably macroscopic systems can exhibit quantum coherence. Unfortunately, Leggett–Garg tests suffer from the ineludible clumsiness loophole—the difficulty of proving that sequential measurements have not inadvertently disturbed the system. The recently uncovered Tsirelson inequality is derived from the simple dynamical assumption of uniform precession, obeyed by many classical systems, and requires only single-time measurements. However, Tsirelson inequality violations could be explained by a macrorealistic system that merely breaks the dynamical assumption, rather than genuine quantum behaviour. By carrying out a quantum-mechanical analysis of the Tsirelson inequality in the harmonic oscillator, we develop a protocol to rule out this possibility by assessing generalised conditions of uniform precession. We show that various measures of uniform precession, some of which related to Leggett–Garg quantities, are satisfied well enough that the presence of quantum-mechanical interference terms must be implied. We derive several incidental mathematical results relating to maximally-violating states, concerning dwell time, crossing number and probability currents, and also consider a group theoretic analysis of the Tsirelson operator.
@@ I.1 The Tsirelson inequality
In 2006, Boris Tsirelson wrote a paper with a title that asks an endearingly straightforward question: “How often is the coordinate of a harmonic oscillator positive?” [1]. Tsirelson demonstrated that a simple dynamical assumption about classical oscillators can give rise to a surprising disparity between classical and quantum expectations. Violations of the “Tsirelson inequality” [2] are associated inherently with the negativity of the Wigner function [3] and present an alluring scheme to assess the “quantumness” of a system using only single-time measurements on an ensemble of systems.
This work went virtually unnoticed for more than a decade and a half until it was rediscovered by Zaw et al., who in their 2022 paper extended Tsirelson’s work to a broader family of systems undergoing uniform precession, such as finite-dimensional spins in real space [4]. Tsirelson’s protocol has been generalised as a viable entanglement witness for coupled harmonic oscillator systems [5, 6] and entangled spin ensembles [7], and the most recent work in the field has focussed primarily upon further generalisations of the protocol [8] and computing stricter bounds on the maximal quantum violations of Tsirelson’s inequality [9, 10]. A recent triumph has been the first experimental realisation (a so-called quantumness certification) of Tsirelson violations in precessing nuclear spins [11].
To derive the Tsirelson inequality, we consider a classical harmonic oscillator with time period $T$. We define the dichotomous variable $Q(t)$ to be the sign function $\text{sign}(x(t))$, which is $\pm 1$ depending on the sign of $x$, and measure this function—which is the mathematical equivalent of asking the question, on which side of the origin does the particle lie?—at three equally-spaced times, every one-third of a period ($t_{1}=0$, $t_{2}=\frac{1}{3}T$, and $t_{3}=\frac{2}{3}T$). A simple dynamical assumption about oscillators is this: the particle will never be found on the same side of the origin three times in a row. Therefore, the sum $Q(0)+Q(\frac{1}{3}T)+Q(\frac{2}{3}T)$ can only assume the values +1+1 or 1-1 and not +3+3, for instance, because this would require $Q=+1$ at all three times. This is illustrated in Fig. 1. (Tsirelson originally considered the harmonic oscillator’s classical phase-space dynamics, illustrated in Fig. 2, and the probability that a particle is found on one side of the origin.)
Figure 1: A classical particle oscillating in a quadratic potential is one example of a system undergoing uniform precession. The dynamical assumption is that the value of $Q$ must take the opposite sign at one of the three measurement times, irrespective of where in its trajectory the particle starts at time $t=0$.
@@
Since the aforementioned dynamical assumption holds true for any system undergoing uniform precession, this Tsirelson inequality would be respected by even, for example, a (classical) charged particle precessing in a magnetic field. However, we focus on the case of the harmonic oscillator for the remainder of this work.
By virtue of the classical equivalence between the expectation value of a sum and the sum of expectation values, $\langle A\rangle$ can just as well be written in terms of single-time averages:
@@
Thus, the Tsirelson quantity $\langle A\rangle$ can be measured using only single-time measurements on an ensemble of identical systems. In practice, measuring the dichotomic variable $Q$ without ascertaining any further information about the particle’s position may be done using a light beam that illuminates one half of the harmonic well [12].
However, a crucial point which has remained hitherto unaddressed in the literature is the validity of the dynamical assumption of uniform precession in a quantum-mechanical context. If truly the Tsirelson test is to be interpreted as a credible certification of quantum behaviour, it is imperative to demonstrate that violations are indicative of genuine quantum phenomena and not simply of the failure of its founding dynamical assumption.
Figure 2: Classical phase-space depiction of the system’s dynamics for three possible measurement times. If the system is initialised in the green region for one of the experimental runs, then it will have a probability of $2/3$ of being found on the right side of the origin at a randomly-chosen measurement time $t\in\{0,T/3,2T/3\}$. Conversely, the probability will be $1/3$ if the particle begins in a blue region. As such, the ensemble average, a convex sum of these probabilities, is bounded between $1/3$ and $2/3$. This is equivalent to Eq. 1.1 up to a scale factor. (Adapted from Zaw et al. [13].)
Uniform precession can be defined perfectly well in the case of a classical oscillator by considering, for instance, the phase-space representation of its trajectory. The condition for uniform precession is then simply that the angular variable $\varphi=\tan^{-1}(x/p)$ evolve linearly in time ($\dot{\varphi}=\text{const.}$). However, quantum-mechanically, the corresponding operator, related to the arrival time operator, is conjugate to the Hamiltonian and is not self-adjoint [14, 15, 16, 17]. As such, the eigenstates of the angular operator cannot be orthogonal, and there is some intrinsic indefiniteness in the angular behaviour and hence in the definition of uniform precession; this property is also related to the difficulties of defining a phase operator for the harmonic oscillator [18]. In practice, this means that uniform precession can be defined in multiple inequivalent ways in quantum mechanics, as we shall see.
The Tsirelson quantity $\langle A\rangle$ may be conveniently written in the Wigner representation [3, 19, 20, 21, 22, 23, 9] in terms of the function $W(p,x,t)$ (and see also Appendix A). It has the form
@@
Although the Wigner function evolves along the classical path, it is not a probability distribution, since it can go negative, and thus we cannot immediately assert that the evolution in the quantum case respects uniform precession. In the Heisenberg picture, with a time-independent Wigner function, the Tsirelson quantity is
@@
We dedicate the remainder of this paper to finding precise sets of conditions for uniform precession in terms of measurable probabilities that are all satisfied classically but which lead to quantum-mechanically inequivalent expressions and measures, which we may use to assess the degree to which the dynamical assumption is satisfied. In order to do so, we will embed the Tsirelson scheme within the more comprehensive Leggett–Garg framework [24]—the canonical test of non-classicality in temporal correlations—and utilise some of its powerful machinery.
I.2 Leggett–Garg tests of non-classicality
@@
The Leggett–Garg inequalities (LGIs) [24, 25], a temporal analogue to Bell-type inequalities [26, 27], must be satisfied by any theory obeying “macrorealism”, a quintessentially classical world-view that is encapsulated by the following three postulates: macroscopic realism per se—a macroscopic system exists in a definite state at all times; non-invasive measurability—the state of the system can be determined, in principle, without affecting its subsequent dynamics; and induction—the current state does not depend upon future measurements.
@@
Macrorealism pertains to measurements made on a specific data set [28], which we take here to be the set of three single-time averages $\langle Q_{i}\rangle$, where $Q_{i}\equiv Q(t_{i})$, and the three correlators $C_{ij}$, the latter measured in a non-invasive way. With possible measurement outcomes $s_{i}=\pm 1$, where $i=1,2,3$ for measurements taken at times $t_{1},t_{2},t_{3}$, respectively, the two-time correlators are defined as
@@
with $1\leq C_{ij}\leq 1$, and where $p_{ij}(s_{i},s_{j})$ is the proportion of outcomes in the data set where a measurement of $Q$ at time $t_{i}$ yields $s_{i}$ and at (a later time) $t_{j}$ yields $s_{j}$.
It will be useful to introduce here some mathematical expressions we will require for later analysis. The measured data set fully determines the two-time quasi-probability distribution [29, 30]
@@
where $D=\langle Q_{1}Q_{2}Q_{3}\rangle$ is the triple correlator but is never measured in the simplest protocols. These two quasi-probabilities are classically positive. Quantum-mechanically, they are equivalent to the expressions
@@
with projection operators $P_{s_{i}}(t_{i})=\frac{1}{2}(1+s_{i}Q_{i})$. We also give here the analogous sequential measurement probabilities and their moment expansions:
@@
where the superscripts denote the presence of an earlier or intermediate measurement that has been summed out. The difference between, for example, $\langle Q_{2}\rangle$ and $\langle Q_{2}^{(1)}\rangle$ signifies the presence of the interference terms which produce the LGI violations.
Turning now to the conditions for macrorealism, the three postulates of macrorealism imply that there exists an underlying joint probability for the given data set, which implies that a set of twelve two-time and four three-time LG inequalities hold. Specifically, the twelve quasi-probabilities Eq. 1.6, for time pairs $(t_{1},t_{2})$, $(t_{1},t_{2})$, $(t_{1},t_{2})$, are non-negative (LG2s) and the following four LG3 inequalities are satisfied:
@@
These sixteen conditions are not only necessary for macrorealism but are also sufficient, as is shown by the LG version of Fine’s theorem [31, 32, 33]. In quantum mechanics, the LG2 and LG3 inequalities can be violated up to the value $\frac{1}{8}$, the Lüders bound. [^1] This is for so-called Lüders measurements, which simply measure the value of $Q$. More negative values can be obtained by measuring finer-grained quantities and then coarse-graining the probabilities to find probabilities of $Q$ [34, 35].
@@
from which we see that the LG3s describe a quasi-probability distribution for changes of sign.
@@
This means that if the LG3s are satisfied but the LG2s are not, then even though MR is violated for the full data set, it can still be satisfied for the data subset consisting of the sign-change variables $\chi_{1}=s_{1}s_{2}$ and $\chi_{2}=s_{2}s_{3}$, which are $\pm 1$ depending on whether or not the sign of $Q$ has flipped between measurement times. This is a point we will make extensive use of in Section III, where we are interested in the sign-change probabilities as a quantification of uniform precession.
@@
Violations of the LGIs have been observed experimentally in a myriad of physical systems, such as superconducting qubits [36, 37], nuclear spins [38, 39, 40], millimetre-scale crystals [41], photons [42], and (recently) neutron interferometry [43, 44]. The Leggett–Garg inequalities also have far-reaching applications to other fields [45, 46, 47, 48, 49]. A more exhaustive account of the experimental tests of macrorealism to date is provided in the extensive review articles Refs. [34, 50].
@@
However, a significant issue is that macrorealism per se is conjoined inextricably with non-invasive measurability in the Leggett–Garg scheme. The “clumsiness loophole” [51] is that any recorded violations—even if a system truly does not admit realism—could always be ascribed, instead, to the unwitting invasiveness of the measurement process. Indeed, a perfectly classical stochastic model with measurements that disturb the system can reproduce correlation functions that violate the LGIs [52]. The standard way to render the measurements non-invasive is to use ideal negative result measurements [53, 54] (and its refinements [55]), but alternative methods exist, such as the waiting detector protocol [56, 57, 58]. If the non-invasive measurement condition is implemented convincingly, then violations of the LGIs imply the failure of macrorealism per se.
@@
The derivation of the Leggett–Garg inequalities solely from the postulates of macrorealism is completely independent of the system’s equations of motion. It is this indifference to the underlying dynamics that necessitates sequential measurements in order to elucidate a quantum–classical disparity. But the prospect of observing non-classicality in temporal correlations need not be plagued by the experimentally challenging task of circumventing invasive sequential measurements. The introduction of a judiciously chosen and experimentally testable dynamical assumption can supplant the requirement for non-invasive measurement. A few examples of such assumptions have appeared in the LG literature and typically consist of assuming that certain transition probabilities are zero [59, 60, 56] or can be simplified in some way, for example via a stationarity assumption [34, 61, 62]. The key point is that the assumptions are chosen so that the correlators simplify to single-time measurement averages. Furthermore, the dynamical assumption is either implemented through the design of the experiment (e.g. the choice of Hamiltonian, where possible) or checked experimentally through separate control experiments. Dynamical assumptions are often classically motivated, and the Tsirelson scheme is an example par excellence, with its simplifying dynamical assumption of uniform precession.
@@
The related single crossing dynamical assumption, studied in the context of a waiting detector model in Ref. [56] and Chapter 6 of Ref. [58], is mathematically equivalent over half a period to the assumption of uniform precession over a full period, an elegant proof of which we provide in Appendix B.
@@
But does a violation truly imply genuine quantum behaviour and the failure of macrorealism, or could a stubborn macrorealist now argue that it stems purely from faulty dynamics? A classical system that breaks the dynamical assumption—that is, one which is not precessing uniformly (or not oscillating at all, even)—will easily violate the Tsirelson inequality. Fig. 3 depicts qualitatively that a Tsirelson violation could be caused by a classical particle that does not obey the expected Newtonian dynamics of oscillation induced by a harmonic potential, lingering for too long on one side of the origin; or perhaps we have just defined the origin inaccurately. Either way, such a trajectory still has a defined position at all times, and hence is macrorealistic, yet a spurious Tsirelson inequality violation will ensue. So, just as separating realism from non–invasive measurability has proved to be the greatest hassle in LG tests, do we now run into a similar problem in Tsirelson tests of separating the failure of realism from this newfound dynamical assumption?
Figure 3: Unlike the blue classical trajectory, which follows the expected sinusoidal dynamics, the erratic red trajectory, still macrorealistic, dwells for too long on the right side of the origin and will violate the Tsirelson inequality, despite there being no quantum behaviour. This is the faulty dynamics loophole.
Our strategy to assess the dynamical assumption will be to re-write the Tsirelson quantity in terms of various probabilities for two-time and three-time measurements that can be interpreted as measures of uniform precession. We will do this in two ways. Our first method involves demonstrating an explicit connection between $\langle A\rangle$ and the probability of zero sign changes, $L_{1}$ (Eqs. 1.12 and 1.16), which should be zero classically. Our second method stems from the interesting observation that a weaker condition of setting merely the difference q(+,+,+)q(,,)q(+,+,+)-q(-,-,-) to be zero also leads to the Tsirelson inequality, as we shall show, and thus may also be considered a measure of uniform precession.
I.3 This paper
Assessing how well the dynamical assumption of uniform precession holds and classifying the nature of Tsirelson violations hereinafter forms the core narrative of this paper. In Section II, we lay forth the mathematical machinery required for quantum-mechanical calculations and discuss quantum states and their violations of the Tsirelson inequality, proving some results relating to maximally-violating states through a group theoretic perspective. We present our most important results in Section III on the main protocol we have devised to assess uniform precession, broken down into the two methods discussed prior. Here we shall require the breadth of the Leggett–Garg formalism to disentangle violations of macrorealism from the breaking of uniform precession. In Section IV, we investigate two other quantities that are less strictly related to the Tsirelson inequality but provide further heuristic evidence of the extent to which uniform precession is satisfied. We will show that one has quantum properties identical to the classical case and the other approximately so. Moreover, we will argue that these results can be used to rule out alternative classical explanations. Finally, we conclude in Section V with a critical assessment of whether the protocol introduced in Section III ultimately sidesteps the problem of non-invasive measurement. We report lengthy technical results and mathematical proofs in the appendices.
II Tsirelson Violations in the Quantum Harmonic Oscillator
II.1 Expansion in energy eigenbasis
Tsirelson showed the unexpected result that probing a quantum harmonic oscillator (QHO) at different times can produce ostensibly non-classical results, despite the fact that the time evolution of quantum operators in the Heisenberg picture abides by the same classical equations of motion.

### v3

@@ Abstract
“Macrorealism” posits that a system possesses definite properties at all times and that we can discover these properties, in principle, without disturbing the system’s subsequent behaviour. The Leggett–Garg inequalities are derived under these assumptions and are readily violated by standard quantum mechanics, thereby providing a scheme to test whether demonstrably macroscopic systems can exhibit quantum coherence. Unfortunately, Leggett–Garg tests suffer from the difficult to avoid clumsiness loophole—the difficulty of proving that sequential measurements have not inadvertently disturbed the system. The recently uncovered Tsirelson inequality is derived from the simple dynamical assumption of uniform precession, obeyed by many classical systems, and requires only single-time measurements. However, Tsirelson inequality violations could be explained by a macrorealistic system that merely breaks the dynamical assumption, rather than genuine quantum behaviour. By carrying out a quantum-mechanical analysis of the Tsirelson inequality in the harmonic oscillator, we develop a protocol to rule out this possibility by assessing generalised conditions of uniform precession. We show that various measures of uniform precession, some of which are related to Leggett–Garg quantities, are satisfied well enough that the presence of quantum-mechanical interference terms must be implied. We derive several incidental mathematical results relating to violating states of Tsirelson’s inequality, concerning dwell time, crossing number and probability currents, and also consider a group theoretic analysis of the Tsirelson operator.
@@ I.1 The Tsirelson inequality
In 2006, Boris Tsirelson wrote a paper with a title that asks an endearingly straightforward question: “How often is the coordinate of a harmonic oscillator positive?” [tsirelson_how_2006]. Tsirelson demonstrated that a simple dynamical assumption about classical oscillators can give rise to a surprising disparity between classical and quantum expectations. Violations of the “Tsirelson inequality” [plavala_tsirelson_2024] are associated inherently with the negativity of the Wigner function [Wigner_function] and present an alluring scheme to assess the “quantumness” of a system using only single-time measurements on an ensemble of systems.
This work went virtually unnoticed for more than a decade and a half until it was rediscovered by Zaw et al., who in their 2022 paper extended Tsirelson’s work to a broader family of systems undergoing uniform precession, such as finite-dimensional spins in real space [precessions]. Tsirelson’s protocol has been generalised as a viable entanglement witness for coupled harmonic oscillator systems [PhysRevLett.130.160201, PhysRevA.108.022211] and entangled spin ensembles [PhysRevA.109.042402]. The most recent work in the field has focussed primarily upon further generalisations of the protocol [PhysRevA.110.062408] and computing stricter bounds on maximal quantum violations [zaw2024threeanglevariantstsirelsonsprecession, PhysRevLett.134.190201]. A recent triumph has been the first experimental realisation (a so-called quantumness certification) of Tsirelson violations in precessing nuclear spins [vaartjes_certifying_2025].
To derive Tsirelson’s inequality, we consider a classical harmonic oscillator with time period $T$. Define the dichotomic variable $Q(t)$ to be the sign function $\text{sign}(x(t))$, which is $\pm 1$ depending on the sign of $x$, and consider its value at three equally-spaced times, every one-third of a period ($t_{1}=0$, $t_{2}=\frac{1}{3}T$, and $t_{3}=\frac{2}{3}T$). A simple dynamical assumption about oscillators is this: the particle will never be found on the same side of the origin three times in a row. Therefore, the sum $Q(0)+Q(\frac{1}{3}T)+Q(\frac{2}{3}T)$ can only assume the values +1+1 or 1-1 for any given classical trajectory. This is illustrated in Fig. 1.
Figure 1: A classical particle oscillating in a quadratic potential is one example of a system undergoing uniform precession. The dynamical assumption is that $Q$ cannot have the same sign at all three times, irrespective of where in its trajectory the particle starts at time $t=0$.
@@
(Tsirelson originally considered the harmonic oscillator’s classical phase-space dynamics, illustrated in Fig. 2, and the probability that a particle is found on one side of the origin.) As the expectation value of a sum equals the sum of expectation values, $\langle A\rangle$ can instead be written in terms of single-time averages:
@@
Thus, the Tsirelson quantity $\langle A\rangle$ can be measured using only single-time measurements on an ensemble of identical systems. Note, in practice, measuring the dichotomic variable $Q$ may be done using a light beam that illuminates one half of the harmonic well [PhysRevLett.120.210402].
The Tsirelson quantity $\langle A\rangle$ may be conveniently written in the Wigner representation [Wigner_function, zachos2005quantum, Hillery1984, Tatarskii1983, Case2008, zaw2024threeanglevariantstsirelsonsprecession] in terms of the function $W(p,x,t)$ (and see also Appendix A). It has the form
@@
In the Heisenberg picture, with a time-independent Wigner function, the Tsirelson quantity is
@@
Figure 2: Classical phase-space depiction of the system’s dynamics for three possible measurement times. If the system is initialised in the green region for one of the experimental runs, then it will have a probability of $2/3$ of being found on the right side of the origin at a randomly-chosen measurement time $t\in\{0,T/3,2T/3\}$. Conversely, the probability will be $1/3$ if the particle begins in a blue region. As such, the ensemble average, a convex sum of these probabilities, is bounded between $1/3$ and $2/3$. This is equivalent to Eq. 1.1 up to a scale factor. (Adapted from Zaw et al. [PhysRevA.106.032222].)
However, a crucial point which has, so far, not been addressed in the literature is the validity of the dynamical assumption of uniform precession in a quantum-mechanical context. If truly the Tsirelson test is to be interpreted as a credible certification of quantum behaviour, it is imperative to demonstrate that violations are indicative of genuine quantum phenomena and not simply of the failure of its founding dynamical assumption.
A classical system that breaks the dynamical assumption—that is, one which is not precessing uniformly (or not oscillating at all, even)—will easily violate the Tsirelson inequality. Fig. 3 depicts qualitatively that a Tsirelson violation could be caused by a classical particle that does not obey the expected Newtonian dynamics of oscillation induced by a harmonic potential, either lingering too long on one side of the origin or oscillating faster than the classical frequency.
Figure 3: Unlike the blue classical trajectory, which follows the expected sinusoidal dynamics, the erratic red trajectories, still macrorealistic, either dwell for too long on one side of the origin or oscillate too rapidly, and will violate the Tsirelson inequality since they are measured on the same side three times in a row.
Although the Wigner function evolves along the classical path, it is not a probability distribution, since it can be negative, and thus we cannot immediately assert that the evolution in the quantum case respects uniform precession. Uniform precession can be defined perfectly well in the case of a classical oscillator by considering, for instance, the phase-space representation of its trajectory. The condition for uniform precession is then simply that the angular variable $\varphi=\tan^{-1}(x/p)$ evolves linearly in time ($\dot{\varphi}=\text{const.}$). However, quantum-mechanically, the corresponding operator, related to the arrival time operator, is conjugate to the Hamiltonian and is not self-adjoint [pauli1958handbuch, Arrival_times, muga2008time, PhysRevA.99.012124]. As such, the eigenstates of the angular operator cannot be orthogonal, and there is some intrinsic indefiniteness in the angular behaviour and hence in the definition of uniform precession; this property is also related to the difficulties of defining a phase operator for the harmonic oscillator [nieto_quantum_1993]. In practice, this means that uniform precession can be defined in multiple inequivalent ways in quantum mechanics, as we shall see.
I.2 This paper
The aim of this paper is to find precise sets of conditions characterising uniform precession in terms of measurable three-time probabilities and then to link them to the Tsirelson inequality. This provides a way to make a clear assessment of the degree to which a Tsirelson inequality violation is caused by the presence of genuine quantum effects or merely explained by a violation of uniform precession.
In order to carry out this aim, we embed the Tsirelson scheme within the more comprehensive Leggett–Garg framework [PhysRevLett.54.857], the canonical test of non-classicality in temporal correlations; this is done in Section II. In Section III, we establish the tools required for quantum-mechanical calculations and discuss quantum violations of the Tsirelson inequality. We present our most important results in Section IV on the main protocol to assess uniform precession, broken down into two methods; both of these utilise the Leggett–Garg formalism. In Section V, we investigate two other quantities—the dwell time and crossing number—that are less strictly related to the Tsirelson inequality but provide further heuristic evidence of the extent to which uniform precession is satisfied. We will show that the former has quantum properties identical to the classical case and the latter approximately so. Moreover, we will argue that these results can be used to rule out alternative classical explanations. Finally, we conclude in Section VI with a critical assessment of whether the protocol introduced in Section IV ultimately sidesteps the problem of non-invasive measurement. We report lengthy technical results and important mathematical proofs in the appendices.
II The Leggett–Garg framework
@@
The Leggett–Garg inequalities (LGIs) [PhysRevLett.54.857, Leggett2008], a temporal analogue to Bell-type inequalities [PhysicsPhysiqueFizika.1.195, PhysRevLett.23.880], must be satisfied by any theory obeying “macrorealism”, a quintessentially classical world-view that is encapsulated by the following three postulates: macroscopic realism per se—a macroscopic system exists in a definite state at all times; non-invasive measurability—the state of the system can be determined, in principle, without affecting its subsequent dynamics; and induction—the current state does not depend upon future measurements.
@@
Macrorealism pertains to measurements made on a specific data set [PhysRevA.103.062212], which we take here to be the set of three single-time averages $\langle Q_{i}\rangle$, where $Q_{i}\equiv Q(t_{i})$, and the three correlators $C_{ij}$, the latter measured in a non-invasive way. With possible measurement outcomes $s_{i}=\pm 1$, where $i=1,2,3$ for measurements taken at times $t_{1},t_{2},t_{3}$, respectively, the three two-time correlators are defined as
@@
with $1\leq C_{ij}\leq 1$ and $\{i,j\}=\{1,2\},\{2,3\},\{1,3\}$, and where $p_{ij}(s_{i},s_{j})$ is the proportion of outcomes in the data set where a measurement of $Q$ at time $t_{i}$ yields $s_{i}$ and at (a later time) $t_{j}$ yields $s_{j}$.
It will be useful to introduce here some mathematical expressions we will require for later analysis. The measured data set fully determines the two-time quasi-probability distribution [PhysRevA.87.022114, klyshko_bell_1996]
@@
where $D=\langle Q_{1}Q_{2}Q_{3}\rangle$ is the triple correlator but is not measured in the simplest protocols. These two quasi-probabilities are classically positive. Quantum-mechanically, they are equivalent to the expressions
@@
with projection operators $P_{s_{i}}(t_{i})=\frac{1}{2}(1+s_{i}\hat{Q}_{i})$ (we avoid using hats for quantum operators except when they are needed to avoid ambiguity between classical and quantum variables). We also give here the analogous sequential measurement probabilities and their moment expansions:
@@
where the superscripts denote the presence of an earlier or intermediate measurement that has been summed out. More specifically, the quantity $\langle Q_{2}^{(1)}\rangle$ is defined as [halliwell_quasi]
$\langle Q_{2}^{(1)}\rangle=\sum_{s_{1},s_{2}}s_{2}p_{12}(s_{1},s_{2})\ ,$ (2.8)
and similarly for $Q_{3}^{(12)}$, $C_{13}^{(2)}$, and $C_{23}^{(1)}$. This is therefore the average of the operator $\sum_{s_{1}}P_{s_{1}}(t_{1}){\hat{Q}}_{2}P_{s_{1}}(t_{1})$. Note that there are two more candidate three-time quasi-probabilities that we will explore, obtained from Eq. 2.3 by replacing $C_{23}$ with $C_{23}^{(1)}$, or replacing $C_{13}$ with $C_{23}^{(2)}$. See Appendix H.
Turning now to the conditions for macrorealism, the three postulates of macrorealism imply that there exists an underlying joint probability for the given data set, which implies that a set of twelve two-time and four three-time LG inequalities hold. Specifically, the twelve quasi-probabilities Eq. 2.2, for time pairs $(t_{1},t_{2})$, $(t_{1},t_{3})$, $(t_{2},t_{3})$, are non-negative (LG2s) and the following four LG3 inequalities are satisfied:
@@
These sixteen conditions are not only necessary for macrorealism but are also sufficient, as is shown by the LG version of Fine’s theorem [halliwell_comparing_2017, halliwell_necessary_2019, PhysRevA.100.042103]. In quantum mechanics, the LG2 and LG3 inequalities can be violated up to the value $\frac{1}{8}$, the Lüders bound. [^1] This is for so-called Lüders measurements, which simply measure the value of $Q$. More negative values can be obtained by measuring finer-grained quantities (von Neumann measurements) and then coarse-graining their probabilities to find probabilities of $Q$ [Emary_2013, budroni2014temporal].
@@
from which we see that, when non-negative, the LG3s describe a probability distribution for changes of sign. Eqs. 2.132.16 are, respectively, the probabilities for zero or two sign changes and the two different possibilities for a single sign change.
@@
This means that if the LG3s are satisfied but the LG2s are not, then even though MR is violated for the full data set, it can still be satisfied for the data subset of the sign-change variables $\chi_{1}=s_{1}s_{2}$ and $\chi_{2}=s_{2}s_{3}$, which are $\pm 1$ depending on whether or not the sign of $Q$ has flipped between measurement times. We will make extensive use of this in Section IV, where we are only interested in the sign-change probabilities as a quantification of uniform precession.
@@
Violations of the LGIs have been observed experimentally in a myriad of physical systems, such as superconducting qubits [palacios-laloy_experimental_2010, PhysRevLett.111.090506], nuclear spins [athalye_investigation_2011, Souza_2011, PhysRevA.87.052102], millimetre-scale crystals [crystals], photons [Goggin_2011], and (recently) neutron interferometry [kreuzgruber_violation_2024, PhysRevA.103.032218]. The Leggett–Garg inequalities also have far-reaching applications to other fields [quantum_computation, quantum_info_processing, PhysRevA.106.012214, wilde_could_2009, Li_2012]. A more exhaustive account of the experimental tests of macrorealism to date is provided in the extensive review articles Refs. [Emary_2013, Vitagliano_2023].
@@
However, a significant issue for the Leggett-Garg scheme is that macrorealism per se is conjoined inextricably with non-invasive measurability. The “clumsiness loophole” [wilde_addressing_2012] is that any recorded violations could always be ascribed, instead, to the unintended invasiveness of the measurement process [montina_dynamics_2012]. The standard way to render the measurements non-invasive is to use ideal negative result measurements [dicke1981interaction, knee2012violation] (and its refinements [PhysRevA.99.022119]), but alternative methods exist, such as the waiting detector protocol [halliwell_leggett-garg_2016, PhysRevA.100.042325, mawbythesis]. If the non-invasive measurement condition is implemented convincingly, then violations of the LGIs imply the failure of macrorealism per se.
@@
The derivation of the Leggett–Garg inequalities solely from the postulates of macrorealism is completely independent of the system’s equations of motion. It is this indifference to the underlying dynamics that necessitates sequential measurements in order to obtain a quantum–classical disparity. Alternatively, the introduction of a well-chosen and experimentally testable dynamical assumption can supplant the requirement for non-invasive measurement. A few examples of such assumptions have appeared in the LG literature and typically consist of assuming that certain transition probabilities are zero [hermens_constraints_2018, LGdecays, halliwell_leggett-garg_2016] or can be simplified in some way [Emary_2013, stationarity2011, PhysRevA.52.R2497]. The key point is that the assumptions are chosen so that the correlators simplify to single-time measurement averages. Furthermore, the dynamical assumption is either implemented through the design of the experiment (e.g. the choice of Hamiltonian, where possible) or checked experimentally through separate control experiments. Dynamical assumptions are often classically motivated, and the Tsirelson scheme is an example par excellence, with its simplifying dynamical assumption of uniform precession.
@@
The related single crossing dynamical assumption, studied in the context of a waiting detector model in Ref. [halliwell_leggett-garg_2016] and Chapter 6 of Ref. [mawbythesis], assumes one crossing of the origin over a chosen time interval and would be valid for a harmonic oscillator over half a period. In fact, this is then mathematically equivalent to the assumption of uniform precession over a full period, a proof of which we provide in Appendix B.
@@
But does a Tsirelson inequality violation truly imply genuine quantum behaviour and the failure of macrorealism, or could a stubborn macrorealist now argue that it stems purely from faulty dynamics? We saw that even trajectories with a defined position at all times can produce spurious Tsirelson inequality violations (Fig. 3). So, just as separating realism from non–invasive measurability has proved to be the greatest hassle in LG tests, do we now run into a similar problem in Tsirelson tests of separating the failure of realism from this newfound dynamical assumption?
Our strategy in Section IV to assess the dynamical assumption will be to use the Leggett–Garg framework to re-write the Tsirelson quantity in terms of various probabilities that can be interpreted as measures of uniform precession.
III Tsirelson Violations in the Quantum Harmonic Oscillator
III.1 Expansion in energy eigenbasis